OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (anonymous):

j = coach johnson's age t = tim's age Coach Johnson is 10 years older than Tim j = t + 10 When Tim started football 3 years ago, Coach Johnson was twice his age t - 3 = (j - 3)/2 And then you have simultaneous equations. j = t + 10 t - 3 = (j - 3)/2 Good luck.

OpenStudy (anonymous):

idk why but i cant solve it.

OpenStudy (anonymous):

Okay, how about use substitution. You know from that first equation what j is; j = t + 10 replace j with (t + 10) in the second equation. so instead of t - 3 = (j - 3)/2 you have t - 3 = ((t + 10) - 3)/2 Now solve for t and you can find out what t is.

OpenStudy (anonymous):

13 is the answer

OpenStudy (anonymous):

Coach is 13 years old

OpenStudy (anonymous):

????????? i get that this a very old question but i have a question for @Coyzita . how is the coach 13 years old ??????